OpenStudy (anonymous):

Just need someone to check my answer. 8x – 12 = 4x + 24 my answer is x=9

OpenStudy (anonymous):

show how you got that in step by step

OpenStudy (lgbasallote):

\[8x - 12 = 4x+ 24\] \[\implies 8x - 4x = 24 + 12\] \[4x = 36\] \[x = 9 \checkmark\] seems right to me

OpenStudy (anonymous):

okay then

OpenStudy (anonymous):

he's the aproover

OpenStudy (anonymous):

yep he is and you are approved lol

OpenStudy (anonymous):

thank you again.

OpenStudy (anonymous):

Yes yours rightt!! :)